# Correcting FLORES Evaluation Dataset for Four African Languages

### v1

@@ Abstract
This paper describes the corrections made to the FLORES evaluation (dev and devtest) dataset for four African languages, namely Hausa, Northern Sotho (Sepedi), Xitsonga and isiZulu. The original dataset, though groundbreaking in its coverage of low-resource languages, exhibited various inconsistencies and inaccuracies in the reviewed languages that could potentially hinder the integrity of the evaluation of downstream tasks in natural language processing (NLP), especially machine translation. Through a meticulous review process by native speakers, several corrections were identified and implemented, improving the dataset’s overall quality and reliability. For each language, we provide a concise summary of the errors encountered and corrected, and also present some statistical analysis that measure the difference between the existing and corrected datasets. We believe that our corrections enhance the linguistic accuracy and reliability of the data and, thereby, contributing to more effective evaluation of NLP tasks involving the four African languages.
@@ 1 Introduction
Low-resource languages, especially from Africa, are greatly under-represented in the Natural Language Processing (NLP) landscape, and this is primarily due to the absence of sufficient resources for both training and evaluation Adelani et al. (2022); Kreutzer et al. (2022). Various efforts are been made to create such resources and these include initiatives by organizations such as Lacuna [^1] https://lacunafund.org/ that fund new and qualitative open datasets, community-driven approaches such as Masakhane [^2] https://www.masakhane.io/, HausaNLP [^3] https://hausanlp.github.io/ and the University of Pretoria’s Data Science for Social Impact (DSFSI) Research Group [^4] https://dsfsi.github.io/, and other individual initiatives Abdulmumin et al. (2022). For machine translation evaluation, the FLORES dataset Goyal et al. (2021); NLLB Team et al. (2022) is widely accepted as a benchmark for evaluation, especially because it was the first of its kind for many languages and enables many-to-many evaluation, making it easier to evaluate say a Hausa to Sepedi translation system without pivoting through a high resource language e.g., English. Recently, the MAFAND dataset Adelani et al. (2022) was created but it only allows bilingual evaluation and is limited to the news domain.
While all these resources are being developed, it is imperative to review them from time to time for validation but also to ensure that resources that are deemed correct and representative are actually are. A revealing work by Kreutzer et al. (2022), albeit on mostly web-crawled datasets, found that a lot of the datasets that are being relied upon for low-resource languages are actually littered with significant errors such as misalignments, incorrect translations, and other issues. The significance of evaluation datasets make them even more deserving of such reviews especially by literate native speakers that know how these languages are used, both in terms of writing and speaking. This paper, therefore, presents a comprehensive review and correction of the public FLORES evaluation datasets for four African languages: Hausa, Northern Sotho, Xitsonga and isiZulu. We also provide the corrected datasets for future evaluation tasks.
@@ 2 The FLORES Evaluation Dataset
This was the original evaluation data and it was created by translating English dataset collected from Wikipedia, consisting of several topics and domains, to 101 mostly low-resource languages. The dataset was the first available evaluation benchmark for several low-resource languages and it enabled the evaluation of many-to-many translation systems without pivoting through another high-resource language such as English. Several quality control mechanisms were put in place to ensure that the final dataset was of acceptable quality. To determine if translations are good enough for inclusion in FLORES-101, a 20% sample of the dataset were reviewed by language-specific evaluators who assess the quality using a Translation Quality Score (TQS) on a 0 to 100 scale, with a score of 90% deemed acceptable. Errors such as grammar, punctuation, spelling, and mistranslation were examined, and each was assigned a severity level of minor, major, or critical. Three of the four languages in this paper were included in this dataset–Hausa (hau), Northern Sotho (nso) and Zulu (zul).
@@ 3.2 Correction Guidelines
lang dev devtest
# corr. # tokenso # tokensc diff. $\Delta$ tokens # corr. # tokenso # tokensc $\Delta$ tokens % div.
hau* 538 15,336 15,239 97 24.3 - - - - -
nso* 62 2,073 2,130 57 29.8 53 1,780 1,819 39 30.5
tso* - - - - - 61 2,127 2,161 34 35.7
zul 192 3,967 3,583 384 25.7 227 4,430 4,423 7 30.4
@@ 4 Error Analysis
A significant part of the translations were incoherent or unclear. Most of them appear to have been generated automatically. As a result, we ran a comparison with the FLORES Hausa translation and new translations by Google Translate. The result showed limited exact matches [^7] Google Translate may have evolved from the time of the dataset creation., notwithstanding, the translation of several words matches with those of Google Translate even when they are obviously wrong. We even found the sentence-level translations generated by Google Translate to be more qualitative and coherent that the present dataset. To illustrate this, we take each sentence from dev and devtest, see Table 3.
@@
Several key challenges and areas for improvement were identified and corrected, focusing on vocabulary consistency, syntax, spelling, and the accurate conveyance of technical terms. Most of the text was accurately translated and, for the text with problems, only small changes were required to make it more accurate.
Some of the words like “safatanaga and disafatanaga” have generally maintained lexical consistency although wrongly translated, which have been corrected to “sefatanaga or difatanaga (plural)”.
Although sometimes Sepedi uses borrowed words for many technical and scientific terms, things such as pavement do have a translation which could be “tsela ya maoto or tselanathoko”. These could have been used instead of borrowing the pavement term to say pabamente. The use of a borrowed term could have been from the available corpus or from learned behaviour for borrowing unknown English terms. Another example is the word college which the model translated to colleje, but Sepedi has an accepted borrowed translation “kholetšhe”.
@@
Some terms were left out, like "scientific" as "tša bo ramahlale" when scientific tools were talked about, and this greatly affected the meaning of the sentence. Additionally, in another instance, a sentence describing the use of Caesarean section to give birth to Nadia was misleading. Incorrectly, it implied that Nadia was both the baby being born and the individual undergoing the operation. This was corrected to have the intended meaning. These errors highlight the importance of precise translations to maintain clarity and accurately convey the original message.
@@
Spelling problems and the usage of borrowed terms can have a substantial influence on the correctness of Xitsonga translations. One of the most illustrative examples of such incongruity of terms is that the English word "channel" has been translated into Macedonian as "chanele". Instead, the work should have used the original term "nongonoko" in order to ensure a perfect linguistic and connotative translation. To avoid generation of wrong impressions, the phrase borrowed from IsiZulu as used to mean "President" had to be replaced by the word "murhangeri wa tiko" from Xitsonga. Deficient spelling, as in the case of writing "dokodela" instead of "Dr", and examples of slang such as using "mwana wa" instead of the formal "muongori" indicate how borrowing and spelling mistakes reduced the quality of the translations. Fluency and correct spelling as well as using the native language correctly are a necessity to maintain the translated material’s effectiveness.
@@
A closer examination of these challenges reveals issues such as in the translation of "Around 11:29, the protest moved up Whitehall, …" which was initially rendered as "Ngawo-11:29 ababhikishi baya Odongeni Olumhlophe, …". This translation contains two key issues. First, "Ngawo-11:29" should have been "Ngabo-11:29" to correctly match the grammatical structure for time expressions in isiZulu. Second, the literal transliteration of "Whitehall" as "Odongeni Olumhlophe" failed to integrate properly into the sentence. The correct approach would involve incorporating the place name with the locative prefix "e-" to produce "e-Whitehall.". This prefix addition is required in conjunctive languages when using borrowed words or terms, but MT systems often fail to capture these variations.
Additionally, another common issue was the unnecessary borrowing of words from English, despite the availability of standardized isiZulu terms. This was particularly evident with month names, scientific terms, and country names, where inconsistencies were frequent—one translation might use "January," another "uJanuwari," and yet another "uMasingana." Another example of this can be seen with the country name "Spain," which was inconsistently translated as both "Spain" and "Speyini" in different sections. Similar inconsistencies occurred with attempts to translate organizational names or acronyms, leading to partial translations that disrupted the linguistic flow.
To address the inconsistencies, standardized isiZulu terms were consistently applied throughout the translations. For instance, month names such as "uMasingana" replaced the inconsistent use of "January" and "uJanuwari." In dealing with organizational names and acronyms and countries’ names, the approach was to fully translate these entities or retain their original form consistently, avoiding partial translations that could disrupt the flow.
@@ 5 Conclusion
In this work, we highlight the importance of qualitative evaluation datasets for low-resource languages and present our findings from a comprehensive review of the FLORES dataset for four African languages: Hausa, Northern Sotho, Xitsonga, and isiZulu. The original translations were marred by vocabulary inconsistencies, syntax errors, and inaccurate technical terms. After making necessary corrections, we measured the amount of edits and resulting difference between the improved datasets and the original using metrics like BLEU, TER, and COMET, which showed that significant improvements were made. Work is still ongoing in some of the languages, as indicated in Table 1, and the final data and its analysis will be included in the camera-ready. Nonetheless, the results presented highlight the need for ongoing refinement and human oversight in developing accurate translation datasets for underrepresented languages. For future work, we intend to expand the corrections to more African languages.
@@ Appendix A Description of the Target Languages
Zulu or isiZulu (in Zulu) is one of the 12 official languages in South Africa, and it is considered to be the most widely spoken indigenous language in the country. It constitutes approximately a quarter of the population, with around 15.1 million speakers out of the population of 62 million people (StatsSA, 2022). IsiZulu is part of the Nguni language family, which is made up of a group of closely related Bantu languages belonging to a larger Niger-Congo language family, and they are widely spoken across Southern Africa (Mesham et al., 2021). These languages are particularly notable for their complex morphology, characterized by agglutinative morphology and conjunctive orthography. Agglutinative morphology means that words are typically formed by combining multiple small meaning-carrying units, known as morpheme. Conjunctive orthography means that the morphemes are glued together to form a word, rather than writing them with spaces in between, as seen in disjunctive orthography, commonly associated with the Sotho group, as well as Tshivenda and Xitsonga in South Africa indigenous languages (Taljard and Bosch, 2006). To illustrate this distinction, consider the example in Figure LABEL:fig:grammatical_structure which examines the different grammatical structures of the phrase I know them.
Table 4 shows that while the phrase’s meaning is consistent across languages, the writing systems vary: in disjunctive orthography, morphemes are separated by spaces, while in conjunctive orthography, as in isiZulu, they are joined into a single word. For example, in the phrase I know them, each morpheme serves a specific grammatical function–’I’ as the subject, ’know’ as the verb, and ’them’ as the object. In disjunctive orthography, these morphemes are written separately, making each unit distinct. In conjunctive orthography, they are combined into one continuous word, but the meaning remains intact. These orthographic variations pose challenges for machine translation systems, which must accurately process morphemes in different writing systems to produce accurate translations.

### v2

@@ Abstract
This paper describes the corrections made to the FLORES evaluation (dev and devtest) dataset for four African languages, namely Hausa, Northern Sotho (Sepedi), Xitsonga, and isiZulu. The original dataset, though groundbreaking in its coverage of low-resource languages, exhibited various inconsistencies and inaccuracies in the reviewed languages that could potentially hinder the integrity of the evaluation of downstream tasks in natural language processing (NLP), especially machine translation. Through a meticulous review process by native speakers, several corrections were identified and implemented, improving the overall quality and reliability of the dataset. For each language, we provide a concise summary of the errors encountered and corrected and also present some statistical analysis that measures the difference between the existing and corrected datasets. We believe that our corrections improve the linguistic accuracy and reliability of the data and, thereby, contribute to a more effective evaluation of NLP tasks involving the four African languages. Finally, we recommend that future translation efforts, particularly in low-resource languages, prioritize the active involvement of native speakers at every stage of the process to ensure linguistic accuracy and cultural relevance.
@@ 1 Introduction
Low-resource languages, especially from Africa, are greatly under-represented in the Natural Language Processing (NLP) landscape, and this is primarily due to the absence of sufficient resources for both training and evaluation Adelani et al. (2022); Kreutzer et al. (2022). Various efforts have been made to create such resources and these include initiatives from organizations such as Lacuna [^1] https://lacunafund.org/ that fund new and qualitative open datasets, and communities such as Masakhane, HausaNLP, the University of Pretoria’s Data Science for Social Impact (DSFSI) Research Group, and other individual initiatives Abdulmumin et al. (2022); Parida et al. (2023). For machine translation evaluation, the FLORES dataset Goyal et al. (2021); NLLB Team et al. (2022) is widely accepted as a benchmark for evaluation, especially because it was the first of its kind for many languages and enables many-to-many evaluation, making it easier to evaluate say a Hausa to Sepedi translation system without pivoting through a high resource language, e.g., English. Recently, the MAFAND dataset Adelani et al. (2022) was created, but it only allows bilingual evaluation and is limited to the news domain.
While all these resources are being developed, it is imperative to review them for validation to ensure that they meet the expected standard of accuracy and representation. A revealing work by Kreutzer et al. (2022), albeit on mostly web-crawled datasets, found that many of the datasets that are being relied upon for low-resource languages are littered with significant errors such as misalignments, incorrect translations, and other issues. The significance of evaluation datasets make them even more deserving of such reviews especially by literate native speakers that know how these languages are written and spoken. This paper, therefore, presents a comprehensive review and correction of the public FLORES evaluation datasets for four African languages: Hausa, Northern Sotho, Xitsonga and isiZulu. We also provide the corrected datasets for future evaluation tasks [^2] https://github.com/dsfsi/flores-fix-4-africa.
@@ 2 The FLORES Evaluation Dataset
This was the original evaluation data and was created by translating the English dataset collected from Wikipedia, consisting of several topics and domains, into 101 mostly low-resource languages. The dataset was the first available evaluation benchmark for several low-resource languages and it enabled the evaluation of many-to-many translation systems without pivoting through another high-resource language such as English. Several quality control mechanisms were put in place to ensure that the final dataset was of acceptable quality. To determine if translations are good enough for inclusion in FLORES-101, a 20% sample of the dataset were reviewed by language-specific evaluators who assess the quality using a Translation Quality Score (TQS) on a 0 to 100 scale, with a score of 90% deemed acceptable. Errors such as grammar, punctuation, spelling, and mistranslation were examined, and each was assigned a severity level of minor, major, or critical. Three of the four languages in this paper were included in this dataset–Hausa (hau), Northern Sotho (nso) and Zulu (zul).
@@
2.1 Problems Identified in FLORES
Prior to this work, we have not found any published work that carefully reviews and attempt to correct mistakes in the FLORES evaluation dataset. However, some issues have been raised on the FLORES’ public GitHub repositories. [^3] https://github.com/openlanguagedata/flores Some of these issues include near-identical translations in several dialects of Arabic: Mesopotamian (acm_arb), Ta’izzi-Adeni (acq_arb), Najdi (ars_arb), and Moroccan (ary_arb) Arabic dialects were found to be too similar to Standard Arabic (arb), [^4] https://github.com/openlanguagedata/flores/issues/8, [^5] https://github.com/facebookresearch/flores/issues/64 unspecifying the "orthography" and "variety" used in Lombard (lmo_latn) and Sardinian (srd_latn), [^6] https://github.com/openlanguagedata/flores/issues/5, [^7] https://github.com/openlanguagedata/flores/issues/6 unmatched quotation marks, [^8] https://github.com/facebookresearch/flores/issues/36 and using Mandarin Chinese in Traditional Chinese Script (zho_Hant) for Cantonese (yue_Hant) translations.
@@ 3.2 Correction Guidelines
lang dev (997 sentences) devtest (1,012 sentences)
# corr. (%) # tokenso # tokensc $\Delta$ tokens % div. # corr. (%) # tokenso # tokensc $\Delta$ tokens % div.
hau 632 (63.4) 17,948 18,073 125 24.7 70 (6.9) 2,006 1,978 28 49.2
nso 67 (6.7) 2,226 2,271 45 28.9 62 (6.1) 2,082 2,105 23 28.0
tso - - - - - 83 (8.2) 2,919 2,947 28 27.4
zul 190 (19.1) 3,605 3,588 17 23.7 226 (22.3) 4,414 4,396 18 31.8
@@
3.3 The Annotators
The correction task was conducted by volunteer annotators that focused on their native languages. These annotators were a mix of university students and researchers holding first, second and third degrees in computing and linguistics.
@@ 4 Error Analysis
A significant part of the translations were suspected to have been automatically generated, as many of them appeared incoherent or unclear. To investigate this, we conducted a comparison with translations from the Hausa FLORES dataset and new translations generated by Google Translate. The comparison revealed that, although there were limited exact matches [^11] Google Translate may have evolved since the creation of the dataset., several incorrect lexical choices in the dataset’s translations aligned with those produced by Google Translate, supporting the suspicion that the translations may have been automatically generated. It is important to note that other translation tools may exist for Hausa that we did not evaluate. Furthermore, several sentence-level translations from Google Translate were found to be more qualitative and coherent than those in the current dataset. For an illustration, we examine sentences from the dev and devtest sets, see Table 3.
@@
Several key challenges and areas for improvement were identified and corrected, focusing on vocabulary consistency, syntax, spelling, and the accurate conveyance of technical terms. Most of the text was accurately translated and, for the text with problems, only small changes were required to make it more accurate. Some of the words like “safatanaga and disafatanaga” have generally maintained lexical consistency although they were wrongly translated. These have been corrected to “sefatanaga or difatanaga (plural)”.
Although sometimes Sepedi uses borrowed words for many technical and scientific terms, things such as pavement do have a translation which could be “tsela ya maoto or tselanathoko”. These could have been used instead of borrowing the pavement term to say pabamente. The use of a borrowed term could have been from the available corpus or from learned behaviour for borrowing unknown English terms. Another example is the word college which was translated to colleje, but Sepedi has a standard borrowed translation: “kholetšhe”.
@@
Some terms were left out, like "scientific" as "tša bo ramahlale" when scientific tools were talked about, and this greatly affected the meaning of the sentence. Additionally, in another instance, a sentence describing the use of Caesarean section to give birth to Nadia was misleading. Incorrectly, it implied that Nadia was both the baby being born and the individual undergoing the operation. This was corrected to have the intended meaning.
@@
Spelling problems and the usage of borrowed terms can have a substantial influence on the correctness of Xitsonga translations. One of the most illustrative examples of such incongruity of terms is that the English word "channel" has been translated as "chanele". Instead, the work should have used the original term "nongonoko" in order to ensure a perfect linguistic and connotative translation. To avoid generation of wrong impressions, the phrase borrowed from IsiZulu as used to mean "President" had to be replaced by the word "murhangeri wa tiko" from Xitsonga. Deficient spelling, as in the case of writing "dokodela" instead of "Dr", and examples of slang such as using "mwana wa" instead of the formal "muongori" indicate how borrowing and spelling mistakes reduced the quality of the translations. Fluency and correct spelling as well as using the native language correctly are a necessity to maintain the translated material’s effectiveness.
@@
A closer examination of these challenges reveals issues such as in the translation of "Around 11:29, the protest moved up Whitehall, …" which was initially rendered as "Ngawo-11:29 ababhikishi baya Odongeni Olumhlophe, …". This translation contains two key issues. First, "Ngawo-11:29" should have been "Ngabo-11:29" to correctly match the grammatical structure for time expressions in isiZulu. Second, the literal transliteration of "Whitehall" as "Odongeni Olumhlophe" failed to integrate properly into the sentence. The correct approach would involve incorporating the place name with the locative prefix "e-" to produce "e-Whitehall.". This prefix addition is required in conjunctive languages when using borrowed words or terms, but MT systems often fail to capture these variations. Additionally, another common issue was the unnecessary borrowing of words from English, despite the availability of standardized isiZulu terms. This was particularly evident with month names, scientific terms, and country names, where inconsistencies were frequent—one translation might use "January," another "uJanuwari," and yet another "uMasingana" Another example of this can be seen with the country name "Spain," which was inconsistently translated as both "Spain" and "Speyini" in different sections. Similar inconsistencies occurred with attempts to translate organizational names or acronyms, leading to partial translations that disrupted the linguistic flow.
To address the inconsistencies, standardized isiZulu terms were consistently applied throughout the translations. For instance, month names such as "uMasingana" replaced the inconsistent use of "January" and "uJanuwari" In dealing with organizational names and acronyms and countries’ names, the approach was to fully translate these entities or retain their original form consistently, avoiding partial translations that could disrupt the flow.
@@ 5 Conclusion
In this work, we highlight the importance of qualitative evaluation datasets for low-resource languages and present our findings from a comprehensive review of the FLORES dataset for four African languages: Hausa, Northern Sotho, Xitsonga, and isiZulu. The original translations were marred by vocabulary inconsistencies, syntax errors, and inaccurate technical terms. After making necessary corrections, we measured the amount of edits and resulting difference between the improved datasets and the original using metrics like BLEU, TER, and COMET, which showed that significant improvements were made. The results presented highlight the need for ongoing refinement and human oversight in developing accurate translation datasets for underrepresented languages. For future work, we intend to expand the corrections to more African languages.
@@ Appendix A Description of the Target Languages
Zulu or isiZulu (in Zulu) is one of the 12 official languages in South Africa, and it is considered to be the most widely spoken indigenous language in the country. It constitutes approximately a quarter of the population, with around 15.1 million speakers out of the population of 62 million people (StatsSA, 2022). IsiZulu is part of the Nguni language family, which is made up of a group of closely related Bantu languages belonging to a larger Niger-Congo language family, and they are widely spoken across Southern Africa (Mesham et al., 2021). These languages are particularly notable for their complex morphology, characterized by agglutinative morphology and conjunctive orthography. Agglutinative morphology means that words are typically formed by combining multiple small meaning-carrying units, known as morpheme. Conjunctive orthography means that the morphemes are glued together to form a word, rather than writing them with spaces in between, as seen in disjunctive orthography, commonly associated with the Sotho group, as well as Tshivenda and Xitsonga in South Africa indigenous languages (Taljard and Bosch, 2006). To illustrate this distinction, consider the example in Table 4 which examines the different grammatical structures of the phrase I know them.
Table 4 shows that while the phrase’s meaning is consistent across languages, the writing systems vary: in disjunctive orthography, morphemes are separated by spaces, while in conjunctive orthography, as in isiZulu, they are joined into a single word. For example, in the phrase I know them, each morpheme serves a specific grammatical function–‘I’ as the subject, ‘know’ as the verb, and ‘them’ as the object. In disjunctive orthography, these morphemes are written separately, making each unit distinct. In conjunctive orthography, they are combined into one continuous word, but the meaning remains intact. These orthographic variations pose challenges for machine translation systems, which must accurately process morphemes in different writing systems to produce accurate translations.